### v1

# What Did I Do Wrong? Quantifying LLMs’ Sensitivity and Consistency to Prompt Engineering

Abstract
Large Language Models (LLMs) changed the way we design and interact with software systems. Their ability to process and extract information from text has drastically improved productivity in a number of routine tasks. Developers that want to include these models in their software stack, however, face a dreadful challenge: debugging their inconsistent behavior across minor variations of the prompt. We therefore introduce two metrics for classification tasks, namely sensitivity and consistency, which are complementary to task performance. First, sensitivity measures changes of predictions across rephrasings of the prompt, and does not require access to ground truth labels. Instead, consistency measures how predictions vary across rephrasings for elements of the same class. We perform an empirical comparison of these metrics on text classification tasks, using them as guideline for understanding failure modes of the LLM. Our hope is that sensitivity and consistency will be powerful allies in automatic prompt engineering frameworks to obtain LLMs that balance robustness with performance.
[…]
What Did I Do Wrong?
Quantifying LLMs’ Sensitivity and Consistency to Prompt Engineering
Federico Errica  and Giuseppe Siracusano  and Davide Sanvito  and Roberto Bifulco NEC Laboratories Europe name.surname@neclab.eu
1 Introduction
There are only two hard things in Computer Science: cache invalidation and naming things. - Phil Karlton
This famous quote refers to the innate tendency of computer scientists to choose poor names for a program’s variables despite the existence of good coding practices (McConnell, 2004). From a purely practical perspective, this is not a problem as long as the program does its job, but can we still argue the same in the era of Large Language Models?
LLMs (Brown et al., 2020) have significantly changed how we process text by providing a straightforward interface, i.e., natural language, to define the problem to be solved (Devlin et al., 2019). They provide software engineers with useful coding tips and can be used as part of larger and more complex software systems (Dakhel et al., 2023). It is common to set up an LLM using a set of instructions called “prompt”, and it soon became clear to both researchers and developers that the prompt itself can greatly influence an LLM’s performance (Zhao et al., 2021; Sclar et al., 2024). The process of writing a good prompt for the current task is called prompt engineering, and a great deal of different techniques have been proposed in this direction (Nori et al., 2023; Sahoo et al., 2024), ranging from a simple description of the problem to providing few-shot examples.
Refer to captionRefer to captionRefer to captionRefer to caption Labels.schema() "What is Australia’s
national flower?"
Function Calling"Description" "Entity" Parser LLM replace ENTY = "An Entity"class Labels(str, enum.Enum):"""Enumeration for single-labeltext classification."""NUM = "Number"DESC = "Description"ENTY = "Entity"
Figure 1: Real example of the behavior of GPT3.5 when classifying a question in terms of what it is referring to. A slight change in the definition of the class “ENTY” causes a minor prompt variation that disrupts the LLM’s prediction. This happens under the hood, making it very hard for a developer to debug the program. Note that the same might happen, for instance, if the ordering of the labels or the fields’ names is changed.
From a practical standpoint, integrating LLMs in a software infrastructure introduces additional complexities, from choosing good prompt engineering strategies to parsing and controlling the output format of responses. In an effort to simplify the LLM into a straightforward function call, new software libraries like Instructor (Liu, 2024) have emerged, reducing even further the entry barrier of LLMs for the ordinary programmer. Instructor requires to define target labels as static fields of a Python class, which is then automatically converted into a standardized JSON schema. The schema and the input sentence are then used to produce an LLM response via a mechanism known as Function Calling (Kang et al., 2023). Figure 1 provides a real example of text classification using Instructor. Here, the assumption of the programmer should be that small changes to the labels’ descriptions, as well as the labels’ ordering in the code, should not affect the final result. Unfortunately, today’s reality is different: a change in the label definition, such as adding an article as shown in the figure, can lead to minor prompt variations with drastic changes in the final prediction. As a result, developers remain unaware of the malfunction’s cause and might later abandon the tool due to frustration.
This simple yet troubling example might be generalized to other problems, for instance, code generation (Liu et al., 2023); at the end of the day, LLM engineers would like to know whether it makes sense to spend their time modifying prompts to obtain a better-behaving LLM; perhaps the LLM does not change its predictions much no matter how the prompt is re-written. Similarly, an LLM whose predictions greatly vary depending on how the prompt is written might generally be regarded as unreliable in a production environment. Therefore, the question we want to address in this paper is the following: "how can we quantify the sensitivity of an LLM to variations of the prompt?". Existing works have answered this question by considering accuracy as the sole metric of interest (McCoy et al., 2023), but this has a limited impact on the everyday life of developers and requires enough ground truth labels for the estimate to be reliable. As a matter of fact, with the recent progress in LLM agents (Gioacchini et al., 2024) and chain of thoughts (Wei et al., 2022) techniques, the existence of multiple intermediate steps and/or user inputs, each handled somehow by an LLM, implies an exponential amount of potential failure paths. LLM engineers need a computationally feasible way to analyze each step individually, possibly irrespective of the final task they need to solve, to reduce the chances that something goes wrong along the way.
We set out to address these problems by proposing two diagnostic metrics for LLM classifiers that are complementary to accuracy: the sensitivity to the input, which does not depend on the ground truth labels, and the consistency of predictions across examples of the same class. Intuitively, it is desirable to have LLMs that are robust to semantically equivalent variations of the initial prompt, and their predicted labels’ distributions should not vary much across samples of the same class. Striving to improve these two metrics might significantly reduce the unpredictability of LLMs’ behavior in complex software systems running in production; low sensitivity and high consistency might mean that LLMs hallucinate less and have a better understanding of the tasks. Consider the example of Figure 2, taken from our experiments, where the predicted labels’ distribution over samples of a given class completely changes for two semantically equivalent but syntactically different rephrasings of the prompt. This behavior is highly undesirable, and the first step to address it is to define suitable metrics to measure it.
Figure 2: Comparison between predicted class distributions of Llama3 across samples of the same class (TREC dataset, Section 4). A merely syntactic rephrasing of the prompt leads to an abrupt change in the distribution, suggesting low consistency.
We evaluate these metrics on five different datasets, two closed and two open-source LLMs, and three prompting strategies, and we find that they indeed convey different information about the LLMs’ behavior. We use this information to analyze more qualitatively some of the datasets, mimicking the kind of study that a user or an LLM engineer would do when debugging their systems. We argue that these metrics should be included in all automatic prompt engineering/optimization frameworks, and perhaps in the training of LLMs themselves, to progress the development of trustworthy AI systems.
2 Related Work
This section positions our work in the context of three different but related research directions: influence of spurious features, uncertainty quantification, and prompt optimization. All these works fall under the broader umbrella of prompt engineering.
Spurious Features.
It is well-known that variations of the prompt affect stability of the LLM’s accuracy. In a broad study across diverse tasks, McCoy et al. (2023) showed that the performance of the LLM depends on the likelihood of the input prompt and the correct output answer. Such result is consistent with the fact that LLMs are, in essence, autoregressive models trained to maximize a likelihood objective. In light of these considerations, it is unsurprising that simply changing the ordering of the examples in a few-shot prompting strategy can lead to almost random accuracy on sentiment analysis tasks (Zhao et al., 2021). Similar considerations motivated frameworks like FormatSpread, which tries to predict the expected performance under prompt’s variations without accessing the LLM’s weights (Sclar et al., 2024). Also, spurious features in the prompt had severe repercussions on the detection of security vulnerabilities, where LLMs are inconsistent and unfaithful (Ullah et al., 2024). Finally, Yang et al. (2024) study the effect of prompt rephrasings and LLMs’ temperature on classification and uncertainty metrics.
Note that the solution to the spurious feature problem might also reside in a more structured approach: Retrieval Augmented Generation systems (Lewis et al., 2020) or Knowledge Graphs-enhanced LLMs (Luo et al., 2024) reduce hallucinations and dependencies on specific prompts.
Uncertainty Quantification.
Complementary to the discussion in this paper is the estimation of the LLM’s uncertainty. Typically, uncertainty is defined over the different answers of the LLM given the same prompt (Press et al., 2024); being generative models of text, LLMs might produce contradicting predictions due to stochasticity in the output response. Several works have already investigated uncertainty: Kadavath et al. (2022) asked the LLM to provide a numerical score of confidence, whereas Chen et al. (2024) argued that aggregating an LLM’s responses via a majority voting mechanism reveals a non-monotonic relationship between the number of LLM calls and system performance. Motivated by the fact that higher uncertainty should imply lower performances, Huang et al. (2024) define a rank calibration error to quantify deviations from the ideal relationship between the two quantities. In a completely different study with about 400 human participants, Kim et al. (2024) provide evidence that allowing an LLM to express uncertainty in the form of natural language reduces the users’ trust in the system. The authors argue that reducing output uncertainty is crucial for LLMs’ successful adoption. Very recently, Yadkori et al. (2024) proposed an information-theoretic metric to distinuish between epistemic uncertainty, which arises from lack of knowledge about the ground truth, and aleatoric uncertainty, which comes from irreducible randomness. They enable the identification of unreliable model outputs and hallucinations without modifying the training process.
While related to our approach, it does not fall within the scope of this paper to further analyze output uncertainty. Because LLMs can also behave deterministically (or close to it) (Yang et al., 2024), which limits the impact of such uncertainty quantification metrics, this work solely focuses on variations of the input prompts. For a comprehensive study of uncertainty quantification techniques on LLMs, we refer the reader to Huang et al. (2023).
Prompt Optimization.
Finally, we discuss techniques to optimize prompts so that the LLMs improve their performance on the task. One cause for bad performances is the bias of the LLM towards samples of a class that appears frequently in the training data; for this reason, a model calibration technique has been proposed to make the predictions more uniform across answers (Zhao et al., 2021). The approach requires access to the LLM’s inner workings, which might not be practical for developers. Automated Prompt Engineer (Zhou et al., 2023) optimizes prompt instructions with the help of an LLM. This approach has also been applied to train the LLM itself, to reduce its susceptibility to adversarial attacks and jailbreaks that overcome the safeguards for ethical use of these systems (Zhou et al., 2024). At the same time, there is still much work to do: LLM-based automatic prompt optimizers struggle to identify the true causes of errors, and we should rather focus on an automated behavior optimization paradigm (Ma et al., 2024).
Prompt optimization is already a step ahead compared to the quantification of sensitivity, and we hope that the definition of suitable metrics will inspire new prompting strategies that do not focus on the sole evaluation of task accuracy.
3 Methodology
We formally introduce the metrics of interest in a bottom-up fashion, namely sensitivity and consistency. We include uncertainty in the discussion to stress the differences between all quantities.
Let us consider a classification task $\tau$ with $C$ possible classes and a data set of $N$ i.i.d. samples $\mathcal{D}={(\bm{x}_{i},y_{i})}_{i\in\{1,\dots,N\}}$, with $\bm{x}$ being the input text to classify and $y\in\{1,\dots,C\}$ the ground truth label. Moreover, we refer to the subset of samples of class $y$ as $\mathcal{D}_{y}$. We make the typical conditional independence assumption about the data, that is, $p(y,\bm{x})=p(y|\bm{x})p(\bm{x})$. Without loss of generality, we introduce the distribution $p_{\rho_{0}}(\rho)$ defined over all prompts that are semantically equivalent to a reference prompt $\rho_{0}$, which is typically prepared by the user based on task $\tau$; when clear from the context, we will omit the subscript from the notation. Because this distribution depends on the task but is independent of the input distribution, we shall write $p(y,\bm{x},\rho)=p(y|\bm{x},\rho)p(\bm{x})p(\rho)$. It is worth noting that LLMs implement the distribution $p(y|\bm{x},\rho)$, hence we can define the average probability that an LLM predicts the class $y$ under different variants of the same prompt as
$p_{\tau}(y|\bm{x})=\mathbb{E}_{\rho\sim p(\cdot)}[p(y|\bm{x},\rho)].$ (1)
In practice, we can approximate Equation 1 by Monte Carlo integration, using $Q$ prompts $\rho_{1},\dots,\rho_{Q}$ semantically equivalent to $\rho_{0}$:
$p_{\tau}(y|\bm{x})\approx\frac{1}{Q}\sum_{i=1}^{Q}p(y|\bm{x},\rho_{i}).$ (2)
We then define the sensitivity of an LLM to the prompt to reflect how much the LLM prediction varies under the rephrasings of the original prompt.
Definition 1 (Sensitivity).
Given a categorical distribution $p_{\tau}(y|\bm{x})$ defined as in Equation 1, the sensitivity with respect to the input $\bm{x}$ is its entropy
$S_{\tau}(\bm{x})=-\mathbb{E}_{y\sim p_{\tau}(\cdot|\bm{x})}[\ln p_{\tau}(y|\bm{x})],$ (3)
whereas the expected sensitivity is
$S_{\tau}=\mathbb{E}_{\bm{x}}[S_{\tau}(\bm{x})]\approx\frac{1}{N}\sum_{i=1}^{N}S_{\tau}(\bm{x}_{i}).$ (4)
It is important to note that the sensitivity does not require access to ground truth labels, which are often hard to acquire, and it does not necessarily correlate with the task’s performance. The sensitivity should be used, for instance, as a guide to compare the “robustness” of different LLMs to variations of the prompt. A highly sensitive LLM may require significant prompt optimization efforts, whereas a less sensitive LLM tells us there might be no further room for improvement. The scope of the prompt variation is also important: one can measure sensitivity w.r.t. minor variations of the original prompt, as well as completely different prompting strategies as long as they convey a semantically equivalent instruction to the LLM. The interpretation we attribute to sensitivity ultimately depends on the use case, but it is universal.
The second metric is called consistency. It measures how much the distribution of Equation 1 differs for two samples $\bm{x},\bm{x}^{\prime}$ of the same class $y$ using the Total Variation Distance (TVD):
$\text{TVD}(p,q)=\frac{1}{2}\sum_{c=1}^{C}|p(c)-q(c)|,$ (5)
whose values range between 00 and $1$.
Definition 2 (Consistency).
Given a categorical distribution $p_{\tau}(y|\bm{x})$ defined as in Equation 1 and two samples $\bm{x},\bm{x}^{\prime}\in\mathcal{D}_{y}$, the pair-wise consistency of a classifier is measured as
$C_{y}(\bm{x},\bm{x}^{\prime})=1-\textnormal{TVD}(p_{\tau}(\cdot|\bm{x}),p_{\tau}(\cdot|\bm{x}^{\prime})],$ (6)
whereas the expected consistency is
$C_{y}=\mathbb{E}_{\bm{x},\bm{x}^{\prime}}[C_{y}(\bm{x},\bm{x}^{\prime})]\approx\sum_{\bm{x},\bm{x}^{\prime}\in\mathcal{D}_{y}}\frac{C_{y}(\bm{x},\bm{x}^{\prime})}{|\mathcal{D}_{y}|^{2}}.$ (7)
Intuitively, a consistent LLM produces similar distributions $p_{\tau}(\cdot|\bm{x})$ regardless of the sample $\bm{x}$ of class $y$, indicating that fine-tuning the LLM on the task $\tau$ might improve the situation. An inconsistent LLM, instead, behaves unpredictably among samples of the same class, suggesting that the main cause of low performance is the prompt $\rho$
We argue that, in order to avoid bad surprises in production environments where new LLMs have to be tested and replaced quickly, it might be desirable to select LLMs with low sensitivity and high consistency, which means $S_{\tau}\rightarrow 0$ and $C_{y}\rightarrow 1,\forall y\in\{1,\dots,C\}$. Jointly optimizing the prompt (or the LLM!) against accuracy, sensitivity, and consistency, might result in less hallucinations and a better understanding of the tasks’s context.
Consider the difference of sensitivity and consistency with respect to the output uncertainty, defined by Huang et al. (2024) as the entropy of the LLM predictions for fixed input $\bm{x}$ and prompt $\rho$:
$U(\bm{x},\rho)=-\mathbb{E}_{y\sim p(\cdot|\bm{x},\rho)}[\ln p(\cdot|\bm{x},\rho)].$ (8)
It is clear that the output uncertainty is orthogonal to sensitivity, and it treats the LLM as a stochastic predictor. However, in classification tasks it is reasonable to force a deterministic behavior of the LLM, effectively achieving $U(\bm{x},\rho)=0\ \forall\bm{x},\rho$. In contrast, as of today, there is no solution for achieving zero sensitivity with an LLM.
4 Experiments
The goal of the experiments is to assess sensitivity and consistency of LLMs on different English-based text classification datasets, analyzing the impact of prompt variations and prompting strategies from a different angle compared to previous works, and showing developers how to use these metrics [^1] https://github.com/nec-research/sensitivity-consistency-LLM.. We ran the LLMs on a 4 Tesla V100 server with 32 GBs of memory and 252 GBs of RAM.
We consider Llama-3-70B-Instruct (Touvron et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024) as the two open-source LLMs available on our servers [^2] Refer to https://huggingface.co/models., as well as GPT3.5-turbo-0125 and GPT-4o-2024-05-13 (Brown et al., 2020) as closed-source models. The temperature is set to zero and the seed is fixed to 42 to obtain quasi-deterministic behavior. Following Zhao et al. (2021), five multiclass classification datasets are used for the comparison: TREC, a 6-class (and 50 subclasses) question-answering task (Voorhees and Tice, 2000) with 500 test samples, CommittmentBank (CB) as a 3-way classification problem (De Marneffe et al., 2019) with 250 test samples, a binary textual entailment problem (RTE, Dagan et al. (2005)) with 2490 test samples, and the 14-class ontology extraction dataset DBPedia (Zhang et al., 2015) with 2000 balanced test samples. In addition, we also consider the 7-class Web of Science 46985 (WoS) dataset with 2000 balanced test samples (Kowsari et al., 2018). In the event the LLM cannot produce a valid class, we add an extra class label N/A.
We consider three different prompting strategies: a simple strategy, where the prompt consists of the task description and the list of classes; a detail strategy, where we provide a detailed description of each class; and a 1-shot strategy in which, compared to simple, we also provide one example for each class taken from samples that do not belong to the test set. To build different rephrasings of the task description only, we use the aforementioned LLMs. In particular, the prompt asks to rephrase the task description by changing the length or adding unnecessary words as long as the meaning remains the same. As done in Yang et al. (2024), we choose $Q=10$. In the quantitative analyses, we display the average sensitivity and average consistency across samples and classes, as well as the micro F1-score across all $\rho$ for completeness. Since the distributions of values are far from being Gaussian, we do not report standard deviations and refer to more qualitative analyses in later sections that show deviations from the mean values.
5 Results
In this section, we want to answer two questions:
i) are sensitivity and consistency complementary to accuracy metrics?; ii) how can we use them to fix prompts, LLMs, and choose the most suitable LLMs for a specific use case?
LLama3 Mixtral
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
TREC .250/.636/.846 .285/.626/.816 .148/.654/.836 .388/.599/.730 .157/.674/.766 .129/.666/.756
CB .025/.639/.932 .020/.637/.924 .008/.644/.916 .245/.563/.704 .005/.640/.932 .176/.652/.944
RTE .149/.590/.721 .214/.591/.764 .047/.588/.818 .182/.581/.780 .165/.584/.739 .041/.587/.827
DBPedia .031/.880/.947 .097/.869/.868 .047/.879/.936 .156/.748/.875 .067/.872/.890 .168/.759/.814
WoS .076/.514/.642 .059/.512/.641 .062/.487/600 .180/.482/.604 .225/.481/.570 .207/.466/.539
GPT3.5 GPT-4o
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
TREC .220/.580/.662 .245/.553/.646 .150/.611/.718 .080/.676/.848 .115/.653/.794 .086/.691/.854
CB .349/.572/.688 .404/.576/.680 .346/.564/.740 .115/.642/.920 .118/.647/.900 .098/.651/.920
RTE .284/.570/.795 .343/.571/.773 .240/.566/.735 .353/.614/.528 .085/.612/.901 .083/.619/.920
DBPedia .040/.873/.955 .074/.864/.906 .047/.870/.949 .061/.886/.914 .063/.879/.925 .053/.878/.954
WoS .200/.515/.632 .190/.517/.619 .196/.511/.653 .097/.515/.658 .093/.515/.658 .090/.511/.660
Table 1: Sensitivity $S_{\tau}$ (lower is better), average consistency $C_{y}$ across all classes (higher is better), and micro-F1 score across all prompt rewritings $\rho$ (higher is better) are shown for different datasets, models, and prompting strategies. Best values across open and closed-source models are shown in bold.
Quantitative Results
To answer question i), Table 1 reports, for each model, dataset, and prompting strategy tried, the values of sensitivity, consistency, and micro F1 score. The first and most important observation is that there seems to be no consistent agreement between the proposed metrics across LLMs and prompting strategies. It is possible that when the F1 score is very high, as in the case of Llama3 with a Simple prompt, the sensitivity is almost double the one achieved with a 1-shot prompt. Similarly, although certainly more related to the task’s performance, the best consistency is rarely associated with the best sensitivity. This result shows that the empirical results agree with our intuitions about the utility of these metrics: a random predictor $p_{\tau}(y|\bm{x})$ would achieve $S_{\tau}\approx 1$ but $C_{y}\approx 1$, whereas another that always predicts a specific class has $S_{\tau}\approx 0$ and $C_{y}\approx 1$. All combinations are possible, and the sensitivity does not depend on the task. As a result, they provide different views about the behavior of the LLM.
Interestingly, the Simple and Detail strategies seem to be more effective across all metrics when using Llama3 and GPT3.5, whereas the Detail and 1-shot ones work better for Mixtral and GPT-4o. This reminds us that developers should always pay attention when switching from one LLM to another in their application scenarios: a prompt that worked well with an LLM might cause instability (see the sensitivity gaps on CB) and significantly worse performance (up to 13% less on DBPedia) on another. For instance, on CB it might be preferable to choose the Detail strategy with Mixtral, since sensitivity is extremely low and the other metrics are close to the best values.
{takeaway}
We observe no consistent pattern of best sensitivity, consistency, and F1; these metrics convey distinct information.
Sensitivity Analysis
(a) TREC
(b) CB
(c) DBPedia
Figure 3: Top: we plot the sensitivity $S_{\tau}$ for each class and prompting strategy (Llama3). Bottom: We show the sensitivity for each sample of the dataset according to different prompting strategies. We remind that the prompting strategy itself might be considered another semantically equivalent rephrasing of the initial prompt $\rho_{0}$ (Section 3).
We move on to answer question ii), by inspecting the sensitivity values against each class and prompting strategy. When some ground truth information is available, Figure 3 (top) can give additional insights into the problematic classes of the task. Consistently with the example of Section 1, we see that on TREC the classes Description, Entity, as well as Number are the ones with the highest sensitivity. This makes sense because the semantic meaning of these two classes can be ambiguous depending on the context. A similar result holds for CB, where the neutral statements are the hardest to classify; here Llama3 is more sensitive to rephrasings of the task description when it comes to neutral statements. This suggests that it might be worth increasing the number of few-shot examples for that class, for instance, or providing a better definition of a neutral statement in the prompt. On DBPedia, instead, it seems that Llama3 has sensitivity issues with the samples of class Artists Building, Animal, and Written Work; as future work, it would be interesting to understand if Llama3 has been trained on less text coming from these domains and relate it to sensitivity.
In Figure 3 (bottom), we visualize the distributions of sensitivity across samples and dataset, divided by prompting strategy. These distributions give us a better intuition of the behavior of the LLM compared to simply checking their mean values. For instance, Llama3 has an average TREC sensitivity of 0.148 with the 1-shot strategy, but there are still a non-negligible number of samples for which sensitivity is still very high. If a developer does not have access to ground truth labels at all, these samples can be manually inspected or given to labelers to identify classes that need more care. As a result, the labeling cost is reduced and the developer knows how to improve the prompt.
{takeaway}
Sensitivity can be used with or without ground truth labels to find “problematic” samples, revealing LLMs’ weak spots.
Consistency Analysis
(a) Person
(b) Entity
(c) Description
Figure 4: Top: we visualize the matrix of pairwise $C_{y}(\bm{x},\bm{x}^{\prime})$ for three different TREC classes, using Llama3 as a classifier. Bottom: we build a histogram for each of the above matrices, to show the distribution of consistency across samples of a given class.
As a reminder, the pair-wise consistency tells us how much the distribution $p_{\tau}(\cdot|\bm{x}),\bm{x}\in\mathcal{D}_{y}$ differs from $p_{\tau}(\cdot|\bm{x}^{\prime}),\bm{x}^{\prime}\in\mathcal{D}_{y}$. Figure 4 (top) provides the matrix $C_{y}(\bm{x},\bm{x}^{\prime})$ for three TREC classes, namely Person, Entity, and Description. Regarding the first class, we consistently observe high consistency except a few cases. By direct observation of the troublesome samples, we can devise prompting strategies targeted to them. In the case of the Entity class, the number of inconsistent pairs is higher than the number of consistent ones, whereby a batch of samples with IDs 46-50 are very consistent with each other but wholly inconsistent against all other samples, e.g., those with IDs 51-58. The former belong to different subclassees (color and other) than the latter (mostly animal), hence defining subclasses in the prompt might provide a better semantic definition of the the class itself. Finally, the matrix of class Description has mixed values, with most of the probability mass being assigned to consistency values smaller than 0.75. Figure 4 (bottom) provides the histogram of these values that convey an aggregated view of these matrices. Compared to Figure 2, it is likely that effort spent improving consistency of the Entity class will also resolve the inconsistencies for the Description class, as these two are often confused by the LLM.
Finally, we analyze the distribution of consistency values across the different prompting strategies, which is shown in Figure 5. There is an interesting pattern on TREC and WoS: the 1-shot strategy reduces the medium-level consistencies, but that does not necessarily imply an increase in all samples’ consistency. There are more pairs of values with consistency 0 compared to not using a Simple or Detail strategy. This is another reason why qualitative results are more helpful than mean values: the behavior of the LLM is counter-intuitive compared to what one would expect, e.g., that a prompting strategy as the few-shot can only increase the consistency. Care should be put when analyzing such behaviors, and the prompt should be adapted accordingly. As regards the CB dataset, the distributions look very similar except for Detail: apparently, providing a class clarification makes the model more sensitive about its predictions for some classes, and we have shown earlier how to debug such faulty behavior.
(a) TREC
(b) DBPedia
(c) WoS
Figure 5: We show the violin plot of the Llama3 consistency over samples of the same classes, arranged by prompting technique, on different datasets.
{takeaway}
Consistency finds sample groups misclassified similarly. Tuning prompts to large groups offers cost-benefit trade-offs.
6 Conclusions
In this work, we have expanded the LLM developer’s toolbox with two diagnostic metrics, namely sensitivity and consistency, that complement performance metrics such as accuracy. In our experiment, we showed how different prompting strategies influence these metrics and how we can guide prompt engineering decisions based on new criteria that value intrinsic characteristics of LLMs’ predictions. Indeed, an LLM that is very sensitive to variation of the prompt and has high test set accuracy might not be a good choice in a production environment, where multiple intermediate steps exist and each could lead to minor alterations in the prompt of the LLM predictor. Notably, sensitivity does not require access to the ground truth labels, and it would be interesting in future work to extend it to tasks different from classification, for instance, code generation. Also, while this work mentioned prompt optimization, it is still an open question how to integrate these metrics in an automatic prompt engineering framework, leading to LLMs insensitive to nonsensical prompt variations and consistent in their (good) performances. As LLM will continue to be neural autoregressive models, and more generally pure neural networks, it is likely that the sensitivity to the input will remain a pain point of these architectures.
7 Limitations
The first clear limitation of the proposed metrics is that they work for classification problems only; despite classification being a common task in information extraction, extending at least sensitivity to more general problems is an important future work. Another inherent limitation is the trade-off between the quality of the approximation used to compute sensitivity and consistency, based on the number $Q$ of different prompt rephrasings, and the cost to run the LLMs. In addition, while the sensitivity can be computed in the absence of ground-truth labels, our analyses revealed that expected values give limited information, and more insights can be gained when we have access to such ground truth, which allows us to understand class-wise sensitivity. Future work should investigate if higher moments, such as the variance, of the metrics we have proposed provide more information without having access to class labels.
8 Ethical Considerations
Our work proposes to evaluate LLMs according to metrics that gauge how well they perform when varying the input prompt and identify failure modes that need solving. For instance, we could use these metrics to discover that an LLM is more sensitive to a minority class than another, allowing us to solve the problem. Malicious attackers can use these metrics to understand if one LLM is more subject to jailbreaks than another, but they give no indication of how to do so. At the same time, making LLMs robust and optimized for these metrics will allow for more trustworthy AI systems that are harder to attack.